# TabEBM: A Tabular Data Augmentation Method with Distinct Class-Specific Energy-Based Models

### v1

@@ Abstract
Data collection is often difficult in critical fields such as medicine, physics, and chemistry. As a result, classification methods usually perform poorly with these small datasets, leading to weak predictive performance. Increasing the training set with additional synthetic data, similar to data augmentation in images, is commonly believed to improve downstream classification performance. However, current tabular generative methods that learn either the joint distribution $p({\mathbf{x}},y)$ or the class-conditional distribution $p({\mathbf{x}}\mid y)$ often overfit on small datasets, resulting in poor-quality synthetic data, usually worsening classification performance compared to using real data alone. To solve these challenges, we introduce TabEBM, a novel class-conditional generative method using Energy-Based Models (EBMs). Unlike existing methods that use a shared model to approximate all class-conditional densities, our key innovation is to create distinct EBM generative models for each class, each modelling its class-specific data distribution individually. This approach creates robust energy landscapes, even in ambiguous class distributions. Our experiments show that TabEBM generates synthetic data with higher quality and better statistical fidelity than existing methods. When used for data augmentation, our synthetic data consistently improves the classification performance across diverse datasets of various sizes, especially small ones.
@@ 1 Introduction
Figure 1: Evaluation of TabEBM and other state-of-the-art tabular generative methods across six key metrics (larger area indicates better performance). The results demonstrate that TabEBM excels in data augmentation (utility), leading to downstream classification improvements by a larger margin than all other methods.
In scientific fields such as medicine, physics, and chemistry, collecting tabular data is often challenging due to the experimental nature of data acquisition [5, 50, 4, 32, 77, 12]. Due to the small size of such datasets [5, 46], training machine learning models that can aid in tasks such as disease diagnosis [52, 37], material property prediction [35], and chemical compound classification [11], often suffer from poor performance [75, 52, 37]. In the case of vision and language tasks, which focus on image and text data, a standard remedy to address model performance due to data scarcity is employing data augmentation techniques [73, 74, 60, 72] that generate additional synthetic samples from existing data. This increases the diversity of the training set and often leads to improved model generalisation [73]. For instance, techniques such as rotating, flipping, or altering the colour of images [73], or generating paraphrased sentences for text [74], have proven effective for many downstream tasks. However, applying data augmentation to tabular data remains understudied, as tabular datasets are very diverse and lack explicit symmetries [8], such as rotations or translations seen in images. Additionally, feature correlations in tabular data are weaker compared to the strong spatial or semantic relationships in image or text data [8]. Consequently, existing tabular data augmentation methods often yield mixed results and can even degrade model performance [51, 72, 48], hindering their widespread adoption.
Tabular augmentation typically involves training generative models to approximate either the joint distribution $p({\mathbf{x}},y)$ [86, 24] or the class-conditional distribution $p({\mathbf{x}}|y)$ [86, 42, 84, 47, 48]. A key challenge of joint distribution methods is maintaining the original training label distribution, as sampling from such generators can produce label distributions that deviate from the original and even fail to generate data for specific classes (see Appendix B for an example). These issues compromise the effectiveness of data augmentation [51] by undermining the label accuracy and distribution. On the other hand, while class-conditional models learning $p({\mathbf{x}}|y)$ preserve the stratification of the original data, they often employ a shared model to represent all class-conditional densities. This, however, can lead to overfitting, particularly in imbalanced datasets where the model may prioritise more frequent classes [21], ignoring unique features needed for generating label-invariant samples. Additionally, in datasets with limited data, this can lead to mode collapse [69, 71], where the model does not effectively capture the diversity of each class [71], and thus tends to perform poorly in a multi-class setting.
@@
  • Library: We release TabEBM as an open-source library (included as a zip file with this submission and to be made publicly available after publication, details in Section A.5). Our library enables off-the-shelf data generation and data augmentation on any tabular dataset without requiring training.
@@ 2.2 Distinct Class-Specific Energy-Based Models
We derive each class-specific EBM $E_{c}({\mathbf{x}})$ by training a classifier on a novel task and reinterpreting its logits. Specifically, for each class $c$, we propose a surrogate binary classification task to determine if a sample belongs to class $c$ by comparing $\mathcal{X}_{c}$ against a set of surrogate negative samples $\mathcal{X}_{c}^{\text{neg}}$, which we show in Figure 3. Specifically, we generate the negative samples at the corners of a hypercube in $R^{D}$. For each dimension $d$, the coordinates of a negative sample are either $\alpha^{\text{neg}}_{\text{dist}}\sigma_{d}$ or $-\alpha^{\text{neg}}_{\text{dist}}\sigma_{d}$, where $\alpha^{\text{neg}}_{\text{dist}}$ is a fixed constant and $\sigma_{d}$ is the standard deviation of dimension $d$. For example, in $R^{3}$, a negative sample might have coordinates $[\alpha^{\text{neg}}_{\text{dist}}\sigma_{1},\alpha^{\text{neg}}_{\text{dist}}\sigma_{2},-\alpha^{\text{neg}}_{\text{dist}}\sigma_{3}]$. Placing the negative samples at the corners of a hypercube ensures they are easily distinguishable from the real data, which is crucial for an accurate energy function (see Section C.1.1). This placement is also robust to variations in the number and distance of the negative samples (see Sections C.1.2 and C.1.3).
@@
where a Gaussian noise term $\epsilon_{t}$ introduces randomness into the sampling process, enhancing the exploration of the distribution. In practice, the step size and the noise standard deviations are often chosen separately, resulting in a biased sampler that allows for faster training. Section C.2 further shows that TabEBM is stable to hyperparameters for the sampling process.
@@
1:Initialise the surrogate binary classification task and train the model
@@ 3 Experiments
Datasets. We utilise 14 open-source tabular datasets–eight from OpenML [7] and six from UCI [22] – across five domains: Medicine, Chemistry, Engineering, Language and Economics. These diverse datasets contain 7 to 77 features and 698 to 5500 samples across 2 to 26 classes. Five datasets contain both numerical and categorical features, while the remaining are numerical only. To provide comprehensive and fine-grained conclusions on model performance, we further enlarge the evaluation scope by varying the degrees of data availability (i.e., $N_{\text{real}}$), leading up to 33 different test cases for the eight OpenML datasets. Section A.1 provide detailed dataset descriptions.
Benchmark generators. We compare TabEBM against eight existing tabular data generation methods of eight different categories: (i) a standard interpolation method SMOTE [13]; (ii) a Variational Autoencoders (VAE) based method TVAE [86]; (iii) a Generative Adversarial Networks (GAN) method CTGAN [86]; (iv) a normalising flow model Neural Spine Flows (NFLOW) [24]; (v) a diffusion model TabDDPM [42]; (vi) a tree-based method Adversarial Random Forests (ARF) [84]; (vii) a Graph Neural Network (GNN) based method GOGGLE [47]; and (viii) a Prior-Data Fitted Networks (PFN) based method TabPFGen [48]. Furthermore, we also include a “Baseline” model, where no data augmentation is applied (i.e., only real data is used to train downstream predictors). In Section A.6, we detail the settings used for TabEBM and all other generators. We have attached the implementation code for TabEBM and will make it publicly available post-publication.
@@
General experimental setup. For each dataset of $N$ samples, we first split it into stratified train and test sets. We create large test sets to reduce the likelihood that the model’s performance is accidentally inflated due to a small, unrepresentative set of samples [70], and thus the test size is computed via $N_{\text{test}}=\min\left(\frac{N}{2},500\right)$. The full train set approximates the upper bound of the quality of synthetic data, and we call this set “oracle”. We subsample the full train set to simulate different levels of data availability, thus the subset size $N_{\text{real}}$ varies over $\{20,50,100,200,500\}$. We split each subset into stratified training and validation sets with a ratio of 4:1. We provide detailed descriptions of data splitting in Section A.2 and preprocessing in Section A.3. We repeat the splitting ten times, summing up to 10 runs per subset size. The reported results are averaged by default over ten runs on the test sets. When aggregating results across datasets, we use the average distance to the minimum (ADTM) metric via affine renormalisation between the top-performing and worse-performing models [30, 54]. We provide the evaluation results averaged over six downstream predictors for a general conclusion, and the fine-grained numerical results for each predictor are in Appendix C.
Data augmentation setup. Given $N_{\text{real}}$ real samples, we first train generators on the real training data and then generate $N_{\text{syn}}$ synthetic samples. For training the downstream predictors, we expand the real training split by adding the synthetic samples. The real validation data is used for early stopping, and the real test set is used for evaluating the predictor’s performance. The optimal $N_{\text{syn}}$ remains an open problem for tabular data [51, 72, 31]. Prior work [47, 48] mainly use synthetic sets with equivalent sizes to the real sets (i.e., $N_{\text{real}}=N_{\text{syn}}$). However, we observe that $N_{\text{real}}=N_{\text{syn}}$ can lead to highly unstable results, especially on small datasets we investigate. Recent work has used different $N_{\text{syn}}$ for various generators, such as by applying post-processing [31, 72]. In this work, we want to provide a head-to-head comparison of the effect of data augmentation across subsampled datasets of varying size $N_{\text{real}}\in\{20,50,100,200,500\}$. Therefore, we perform data augmentation with a large synthetic set ($N_{\text{syn}}=500$) across all splits, and the synthetic data has the same class distribution as the real training data. We provide an illustrative figure of the data splitting setup in Section A.2.
@@ 3.1 Data Augmentation Improvement (Q1)
TabEBM effectively improves downstream performance across sample sizes, especially for very low-sample-size regimes. Table 1 and Figure 4 (Left) show that TabEBM exhibits competitive performance in data augmentation, generally achieving the highest downstream accuracy and average rank across most datasets and sample sizes. Notably, TabEBM is the only generator that consistently improves performance across sample sizes. A key observation is that most modern benchmark generators even underperform the Baseline, indicating poor approximated distributions in the low-sample-size regime. Moreover, TabEBM achieves the largest overall performance improvement on six additional UCI datasets, further supporting its effectiveness (see Section C.5.2 for detailed results).
@@ 3.2 Statistical Fidelity (Q2)
We evaluate the fidelity of synthetic data by focusing on two aspects of synthetic data: similarity to real train data and real test data. We evaluate the similarity between real and synthetic data via (i) average inverse of the Kullback–Leibler Divergence (inverse KL) [17], (ii) p-value of Kolmogorov–Smirnov test (KS test) [39] and (iii) p-value of Chi-squared test ($\chi^{2}$ test) [55]. The full results, including $\chi^{2}$ test and the similarity between real test data and synthetic data, are in Section C.6. For all three metrics, a bigger value denotes that synthetic data is more likely to have the same distribution as real data.
In Figure 7 (a1&a2), TabEBM consistently achieves the highest accuracy and distribution similarity between real train data and synthetic data, showing that TabEBM learns the distributions of real train data better than benchmark generators. In Section C.6, we further show that TabEBM remains the most competitive method in similarity between real test data and synthetic data. This indicates that TabEBM can extrapolate beyond real train data and thus generate synthetic data from a more general distribution that aligns with both train and test data. This extrapolation ability also explains why TabEBM can outperform Baseline via data augmentation (Section 3.1).
@@ 3.3 Privacy Preservation (Q3)
Specifically, we evaluate synthetic data via three metrics: (i) balanced accuracy of downstream predictors trained with only synthetic data (i.e., train-on-synthetic, test-on-real [86, 42, 88]); (ii) median Distance to Closest Record (DCR) [89], where a greater DCR denotes synthetic data is less likely to be copied from real data; and (iii) ${\delta}$-presense [62], where a smaller value denotes a lower re-identification risk for real data from synthetic data. Full numerical results are in Section C.7.
@@
Figure 7: (a1&a2): Median inverse KL and KS test vs. mean normalised balanced accuracy improvement (%) between real train data and synthetic data. (b1&b2): Median DCR and $\delta$-presence vs. mean normalised balanced accuracy change (%) between real train data and synthetic data. Note that “accuracy improvement” is for data augmentation, and “accuracy change” is for data sharing. The figures with with standard deviations are in Section C.4. TabEBM generates high-fidelity synthetic data that can also be used for privacy preservation.
@@ 3.4 Why is TabEBM effective for estimating Energy-Based Models? (Q4)
We found it essential to place the negative samples far from the real data, since TabPFN, which pre-trained to approximate Bayesian inference [33], has its confidence influenced by the distance from the training data [53]. Figure 8 (left) shows that TabPFN outputs high logit values near the real data. As the distance from the real data increases, the logit $f({\mathbf{x}})[1]$ decreases smoothly until the two logits become similar, making the classifier uncertain (because the class probabilities become equal). The right figure shows that TabEBM’s inferred density drops significantly as the maximum logit decreases, because $p_{c}({\mathbf{x}})\propto(\exp(f({\mathbf{x}})[0])+\exp(f({\mathbf{x}})[1]))$ from Equation 3. Since SGLD sampling performs gradient ascent on the density, the TabEBM-generated samples will be close to the real data. These findings are consistent across datasets (see Section C.3), where TabPFN’s logits remain positive, with similar ranges and a relatively constant sum as distance increases, warranting further investigation. Overall, TabPFN’s distance-based uncertainty is useful for inferring accurate energy functions within our TabEBM framework. Since TabEBM can be paired with any other gradient-based classifier that produces logits, we leave these extensions for future work.

### v2

@@ Abstract
Data collection is often difficult in critical fields such as medicine, physics, and chemistry. As a result, classification methods usually perform poorly with these small datasets, leading to weak predictive performance. Increasing the training set with additional synthetic data, similar to data augmentation in images, is commonly believed to improve downstream classification performance. However, current tabular generative methods that learn either the joint distribution $p(\mathbf{x},y)$ or the class-conditional distribution $p(\mathbf{x}\mid y)$ often overfit on small datasets, resulting in poor-quality synthetic data, usually worsening classification performance compared to using real data alone. To solve these challenges, we introduce TabEBM, a novel class-conditional generative method using Energy-Based Models (EBMs). Unlike existing methods that use a shared model to approximate all class-conditional densities, our key innovation is to create distinct EBM generative models for each class, each modelling its class-specific data distribution individually. This approach creates robust energy landscapes, even in ambiguous class distributions. Our experiments show that TabEBM generates synthetic data with higher quality and better statistical fidelity than existing methods. When used for data augmentation, our synthetic data consistently improves the classification performance across diverse datasets of various sizes, especially small ones. Code is available at https://github.com/andreimargeloiu/TabEBM.
@@ 1 Introduction
Figure 1: Evaluation of TabEBM and other state-of-the-art tabular generative methods across six key metrics (larger area indicates better performance). The results demonstrate that TabEBM excels in data augmentation (utility), with a larger area than all other methods.
Many scientific domains within medicine, physics, and chemistry often rely on intricate and challenging data acquisition procedures [5, 50, 4, 32, 76, 12], that typically render small-size tabular datasets [5, 46]. Using these to train machine learning models, that can aid in tasks such as disease diagnosis [52, 37], material property prediction [35], and chemical compound classification [11], can lead to poor performance [74, 52, 37]. In the case of learning tasks which leverage image and text data, a standard remedy to address performance issues due to data scarcity is employing data augmentation techniques [72, 73, 60, 71] that generate additional synthetic samples from existing data. However, applying data augmentation to tabular data introduces additional challenges, as tabular datasets are often very diverse and lack explicit symmetries [8]. Consequently, existing tabular data augmentation methods often yield mixed results and can even degrade model performance [51, 71, 48], hindering their widespread adoption.
Tabular augmentation typically involves training generative models to approximate either the joint distribution $p({\mathbf{x}},y)$ [85, 24] or the class-conditional distribution $p({\mathbf{x}}|y)$ [85, 42, 83, 47, 48]. A key challenge of joint distribution methods is maintaining the original training label distribution, as sampling from such generators can produce label distributions that deviate from the original and even fail to generate data for specific classes (see Appendix C for an example). These issues compromise the effectiveness of data augmentation [51] by undermining the label accuracy and distribution. On the other hand, while class-conditional models learning $p({\mathbf{x}}|y)$ preserve the stratification of the original data, they often employ a shared model to represent all class-conditional densities. This, however, can lead to overfitting, particularly in imbalanced datasets where the model may prioritise more frequent classes [21], ignoring unique features needed for generating label-invariant samples. Additionally, in datasets with limited data, this can lead to mode collapse [68, 70], where the model does not effectively capture the diversity of each class [70], and thus tends to perform poorly in a multi-class setting.
@@
  • Library: We release TabEBM as an open-source library, available at https://github.com/andreimargeloiu/TabEBM. Our library enables off-the-shelf data generation and data augmentation on any tabular dataset without requiring training. Further details are available in Section B.5.
@@ 2.2 Distinct Class-Specific Energy-Based Models
We derive each class-specific EBM $E_{c}({\mathbf{x}})$ by training a classifier on a novel task and reinterpreting its logits. Specifically, for each class $c$, we propose a surrogate binary classification task to determine if a sample belongs to class $c$ by comparing $\mathcal{X}_{c}$ against a set of surrogate negative samples $\mathcal{X}_{c}^{\text{neg}}$, which we show in Figure 3. Specifically, we generate the negative samples at the corners of a hypercube in $R^{D}$. For each dimension $d$, the coordinates of a negative sample are either $\alpha^{\text{neg}}_{\text{dist}}\sigma_{d}$ or $-\alpha^{\text{neg}}_{\text{dist}}\sigma_{d}$, where $\alpha^{\text{neg}}_{\text{dist}}$ is a fixed constant and $\sigma_{d}$ is the standard deviation of dimension $d$. For example, in $R^{3}$, a negative sample might have coordinates $[\alpha^{\text{neg}}_{\text{dist}}\sigma_{1},\alpha^{\text{neg}}_{\text{dist}}\sigma_{2},-\alpha^{\text{neg}}_{\text{dist}}\sigma_{3}]$. Placing the negative samples at the corners of a hypercube ensures they are easily distinguishable from the real data, which is crucial for an accurate energy function (see Section D.1.1). This placement is also robust to variations in the number and distance of the negative samples (see Sections D.1.2 and D.1.3).
@@
where a Gaussian noise term $\epsilon_{t}$ introduces randomness into the sampling process, enhancing the exploration of the distribution. In practice, the step size and the noise standard deviations are often chosen separately, resulting in a biased sampler that allows for faster training. Section D.2 further shows that TabEBM is stable to hyperparameters for the sampling process.
@@
1:Initialise a surrogate binary classification task and train the model
@@ 3 Experiments
Datasets. We utilise eight open-source datasets from OpenML [7] – across five domains: Medicine, Chemistry, Engineering, Language and Economics. We note that TabPFN almost drains the small-size OpenML datasets for its meta-validation [33], which might lead to data leakage when evaluating TabEBM. To provide fair comparisons, we further select six leakage-free datasets from UCI [22]. These diverse datasets contain 7 to 77 features and 698 to 5500 samples across 2 to 26 classes. Five datasets contain both numerical and categorical features, while the remaining are numerical only. We further enlarge the evaluation scope by varying the degrees of data availability (i.e., $N_{\text{real}}$), leading up to 33 different test cases for the eight OpenML datasets. Section B.1 provide detailed descriptions.
Benchmark generators. We compare TabEBM against eight existing tabular data generation methods of eight different categories: (i) a standard interpolation method SMOTE [13]; (ii) a Variational Autoencoders (VAE) based method TVAE [85]; (iii) a Generative Adversarial Networks (GAN) method CTGAN [85]; (iv) a normalising flow model Neural Spine Flows (NFLOW) [24]; (v) a diffusion model TabDDPM [42]; (vi) a tree-based method Adversarial Random Forests (ARF) [83]; (vii) a Graph Neural Network (GNN) based method GOGGLE [47]; and (viii) a Prior-Data Fitted Networks (PFN) based method TabPFGen [48]. Furthermore, we also include a “Baseline” model, where no data augmentation is applied (i.e., only real data is used to train downstream predictors). In Section B.6, we detail the settings used for TabEBM and all other generators.
@@
General experimental setup. For each dataset of $N$ samples, we first split it into stratified train and test sets. We create large test sets to reduce the likelihood that the model’s performance is accidentally inflated due to a small, unrepresentative set of samples [69], and thus the test size is computed via $N_{\text{test}}=\min\left(\frac{N}{2},500\right)$. The full train set approximates the upper bound of the quality of synthetic data, and we call this set “oracle”. We subsample the full train set to simulate different levels of data availability, thus the subset size $N_{\text{real}}$ varies over $\{20,50,100,200,500\}$. We split each subset into stratified training and validation sets with a ratio of 4:1. We provide detailed descriptions of data splitting in Section B.2 and preprocessing in Section B.3. We repeat the splitting ten times, summing up to 10 runs per subset size. The reported results are averaged by default over ten runs on the test sets. When aggregating results across datasets, we use the average distance to the minimum (ADTM) metric via affine renormalisation between the top-performing and worse-performing models [30, 54]. We provide the evaluation results averaged over six downstream predictors for a general conclusion, and the fine-grained numerical results for each predictor are in Appendix D.
Data augmentation setup. Given $N_{\text{real}}$ real samples, we first train generators on the real training data and then generate $N_{\text{syn}}$ synthetic samples. For training the downstream predictors, we expand the real training split by adding the synthetic samples. The real validation data is used for early stopping, and the real test set is used for evaluating the predictor’s performance. The optimal $N_{\text{syn}}$ remains an open problem for tabular data [51, 71, 31]. Prior work [47, 48] mainly use synthetic sets with equivalent sizes to the real sets (i.e., $N_{\text{real}}=N_{\text{syn}}$). However, we observe that $N_{\text{real}}=N_{\text{syn}}$ can lead to highly unstable results, especially on small datasets we investigate. Recent work has used different $N_{\text{syn}}$ for various generators, such as by applying post-processing [31, 71]. In this work, we want to provide a head-to-head comparison of the effect of data augmentation across subsampled datasets of varying size $N_{\text{real}}\in\{20,50,100,200,500\}$. Therefore, we perform data augmentation with a large synthetic set ($N_{\text{syn}}=500$) across all splits, and the synthetic data has the same class distribution as the real training data. We provide an illustrative figure of the data splitting setup in Section B.2.
@@ 3.1 Data Augmentation Improvement (Q1)
TabEBM effectively improves downstream performance across sample sizes, especially for very low-sample-size regimes. Table 1 and Figure 4 (Left) show that TabEBM exhibits competitive performance in data augmentation, generally achieving the highest downstream accuracy and average rank across most datasets and sample sizes. Notably, TabEBM is the only generator that consistently improves performance across sample sizes. A key observation is that most modern benchmark generators even underperform the Baseline, indicating poor approximated distributions in the low-sample-size regime. Moreover, TabEBM achieves the largest overall performance improvement on six leakage-free UCI datasets, further supporting its effectiveness (see Section D.5.2 for details).
@@ 3.2 Statistical Fidelity (Q2)
We evaluate the fidelity of synthetic data by focusing on two aspects of synthetic data: similarity to real train data and real test data. We evaluate the similarity between real and synthetic data via (i) average inverse of the Kullback–Leibler Divergence (inverse KL) [17], (ii) p-value of Kolmogorov–Smirnov test (KS test) [39] and (iii) p-value of Chi-squared test ($\chi^{2}$ test) [55]. The full results, including $\chi^{2}$ test and the similarity between real test data and synthetic data, are in Section D.6. For all three metrics, a bigger value denotes that synthetic data is more likely to have the same distribution as real data.
In Figure 7 (a1&a2), TabEBM consistently achieves the highest accuracy and distribution similarity between real train data and synthetic data, showing that TabEBM learns the distributions of real train data better than benchmark generators. In Section D.6, we further show that TabEBM remains the most competitive method in similarity between real test data and synthetic data. This indicates that TabEBM can extrapolate beyond real train data and thus generate synthetic data from a more general distribution that aligns with both train and test data. This extrapolation ability also explains why TabEBM can outperform Baseline via data augmentation (Section 3.1).
@@ 3.3 Privacy Preservation (Q3)
Specifically, we evaluate synthetic data via three metrics: (i) balanced accuracy of downstream predictors trained with only synthetic data (i.e., train-on-synthetic, test-on-real [85, 42, 87]); (ii) median Distance to Closest Record (DCR) [88], where a greater DCR denotes synthetic data is less likely to be copied from real data; and (iii) ${\delta}$-presense [62], where a smaller value denotes a lower re-identification risk for real data from synthetic data. Full numerical results are in Section D.7.
@@
Figure 7: (a1&a2): Median inverse KL and KS test vs. mean normalised balanced accuracy improvement (%) between real train data and synthetic data. (b1&b2): Median DCR and $\delta$-presence vs. mean normalised balanced accuracy change (%) between real train data and synthetic data. Note that “accuracy improvement” is for data augmentation, and “accuracy change” is for data sharing. The figures with with standard deviations are in Section D.4. TabEBM generates high-fidelity synthetic data that can also be used for privacy preservation.
@@ 3.4 Why is TabEBM effective for estimating Energy-Based Models? (Q4)
We found it essential to place the negative samples far from the real data, since TabPFN, which pre-trained to approximate Bayesian inference [33], has its confidence influenced by the distance from the training data [53]. Figure 8 (left) shows that TabPFN outputs high logit values near the real data. As the distance from the real data increases, the logit $f({\mathbf{x}})[1]$ decreases smoothly until the two logits become similar, making the classifier uncertain (because the class probabilities become equal). The right figure shows that TabEBM’s inferred density drops significantly as the maximum logit decreases, because $p_{c}({\mathbf{x}})\propto(\exp(f({\mathbf{x}})[0])+\exp(f({\mathbf{x}})[1]))$ from Equation 3. Since SGLD sampling performs gradient ascent on the density, the TabEBM-generated samples will be close to the real data. These findings are consistent across datasets (see Section D.3), where TabPFN’s logits remain positive, with similar ranges and a relatively constant sum as distance increases, warranting further investigation. Overall, TabPFN’s distance-based uncertainty is useful for inferring accurate energy functions within our TabEBM framework. Since TabEBM can be paired with any other gradient-based classifier that produces logits, we leave these extensions for future work.